The z-score is used to convert a variable with a Gaussian [Normal] distribution with mean m and standard error s to a variable with a standard normal distribution. Since the latter is tabulated, the probability of an outcome as extreme or more compared to the one observed is easily obtained.
Why standard deviation is better measure of variance?
1. Standard deviation is not a measure of variance: it is the square root of the variance.
2. The answer depends on better than WHAT!
What is the probability of tossing at least one head with 3 coins?
3 coins can land in 8 different ways. Only one of these ways is all tails.
So the probability of rolling at least one heads is 7/8 = 87.5% .
What is the probability of rolling a 4 each time when a 6 sided die is rolled three times?
1/6 x 1/6 x 1/6 = 1/216
You can color ANY map on a flat surface, or on a sphere, using only four colors. With the large number of states which the United States has (48 in continental United States), I strongly suspect that you can't do it with less than that.
soory dude or girl but this was just a joke question. but i hope you find what your looking for.
by Michael Johnson
What are some websites that give advice and have videos about becoming a good soccer player?
i dont no, but i give GREAT advice! ( but sometimes it doen't work out.. but that is VERY rare 4 me! :0) I think if u want 2 b a better soccer player, u have 2 practice as much as u can. u have 2 ask and pray 2 God & ask Him 2 help u b a better soccer player! i hope i hav helped u!! bye! ( p.s. & remember... stay in skool & ALWAYS REMEMBER THIS: Jesus loves u soooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooooo much!!) Click on the 'Soccer Training Tips' link on this page for advice on becoming a better soccer player.
What is the Distance of 22 short?
Are you asking about the maximum effective range or the maximum range that a 22 short projectile can travel?
How many cards can you have in a magic deck?
you can have any number above 60 cards in a deck you just have to be able to shuffle it with out assistance
(1/2)*(2/6) = 1/6
1/6 on the first roll, and 1/6 of rolling a six on the second roll. Overall, the chances of you rolling a 5 on the first and a six on the second is 1/3.
Probability of drawing a 10 out of 52 cards?
The probability of drawing a 10 out of 52 cards is 4 in 52, or 1 in 13, or about 0.07692.
When studying the behaviour of the sum of independent, identically distributed (iid) Gaussian (normal) variables. The F-distribution is used mainly in the analysis of variance where the errors - between the observed value and those predicted by a model - are assumed to be iid Normal.
There are 26 possible letters and 10 possible numbers. The number of license plates possible is then 26*26*10*10*10*10 = 6760000.
How do you find the theoretical probability of an outcome?
The first step is to find a model that adequately represents the situation. You then apply reasoning based on the laws of science, along with some assumption regarding the model, to find out how likely a given outcome is. That value is its theoretical probability.
What is the probability of flipping no heads when you flip three identical coins?
The probability of flipping one coin and getting tails is 1/2. In order to find the probability of multiple events occurring, you find the product of all the events. For 3 coins the probability of getting tails 3 times is 1/8 because .5 x .5 x .5 = .125 or 1/8.
What does it mean to find the probability of an event?
Read the introduction to probability and probability measures at StatLect.com
What is the probability of rolling a sum of 9 or less Suppose you roll two dice one time only?
It is 5/6.
It is necessary to have an estimate of the proportion of families in the town which own cars. Since the number of families which do own a car is given, it is necessary to know the total number of families in the town. Without the total number of families, it is not possible to estimate the car-owning proportion and therefore impossible to make any estimate about the answer.
If, for example, if there were only 65 families this would imply that every family owned a car and, therefore, the probability of exactly 5 out of 7 random families owning a car would be 0. If there were 67 families, the probability would be approx 0.17 while if there were 270 or more families, the probability is zero.
What is the probability that a year selected at random is a leap year with explanation?
The probability is very close to 0.25
A year is a leap year
if the number is divisible by 4 - except
if the number is divisible by 100 it is not a leap year - except
if the number is divisible by 400 it is a leap year.
So, in a 400-year period there are 97 leap years. The probability or relative frequency of leap years is, therefore, 97/400 = 0.2425
What can be probability having 5 Sunday in June?
1 8 15 22 29 or
2 9 16 23 30 so
only if 1st or 2nd is a Sunday will there be five Sundays in the 30-day moth of June. As there are seven days in a week then the odds are 2 in 7